OpenStudy (anonymous):

Find the exact value of the following function. Show your work and explain the steps you take to find the value. Sec (17π/6)

OpenStudy (anonymous):

\[\sec (17\pi /6)\] = \[\sec ((12\pi + 3\pi + 2\pi)/6)\] =\[\sec (2\pi + \pi/2 + \pi/3)\] =\[\sec (\pi/2 + \pi/3)\] =\[\csc (\pi/3)\] = \[2/\sqrt{3}\] =1.732